A 4 digit code number is generated by randomly selecting digits, with replacement, from the set {1,2,3,...,9}. Find the probability that the number consists of two 9's and two 6's.Give an exact answer as a fraction of the form a/b (explained).
Answer #1

total number of possible 4 digit code with replacement from 9 digits =94 =6561

number of ways 4 digit code can be made with two 9's and two 6's =4!/(2!*2!) =6

therefore probability that the number consists of two 9's and two 6's =6/6561 =2/2187